# Forgotten Polygons: Multimodal Large Language Models are Shape-Blind

### v3

@@ Abstract
William Rudman⋆1, Michal Golovanevsky⋆1, Amir Bar2, Vedant Palit3, Yann LeCun4, Carsten Eickhoff5, Ritambhara Singh1 1Brown University, 2Tel Aviv University, 3IIT Kharagpur, 4New York University, 5University of Tübingen {william_rudman, michal_golovanesky}@brown.edu
@@ 3.1.1 What does the text-decoder know?
To assess LLMs’ knowledge of the geometric properties MLLMs struggle with, we construct the following prompts: (1) “What is the name of a <n>-sided polygon?” where <n> ranges from 3 to 8, and (2) “How many sides does a <shape> have?” where <shape> is a regular polygon from triangle to octagon. The results, shown in Appendix Table 7, reveal that the underlying language models know geometric properties. All models get 100% accuracy on “How many sides does a <shape> have?”, and 11 out of 13 models get 100% accuracy on “What is the name of a <n>-sided polygon?”. The success of LLM backbones in these tasks aligns with prior findings that state-of-the-art models perform well on lower-level math problems Xiang et al. (2025). Note that LLaVA 1.5 and G-LLaVA’s LMs struggle with naming certain shapes, suggesting some failures stem from gaps in pretraining data rather than vision-related issues. In Appendix E, we analyze Google N-Grams data and confirm that the shapes these models struggle with are also the least frequently mentioned in text.
@@ 3.3 Counting Sides of Abstract Shapes
Merged Shapes: We combine two regular polygons, such as a triangle and a square, where the shapes share a side. A simple shared-side algorithm can be applied to obtain the number of the sides: $s_{1}+s_{2}-2$ where $s_{1}$ and $s_{2}$ are the number of sides of shape 1 and shape 2, respectively. Applying this algorithm to a triangle and a square gives: $3+4-2=5$ total sides. Table 3 shows that models generally fail to implement this algorithm, with most achieving 0% accuracy on merged shapes (Tri-Sq through Sq-Hep). LLaVA-OneVision and LLaMA-3.2 each achieve over 90% accuracy on exactly one merged shape. This is because most open-source models (except for Molmo, which is trained on counting) tend to guess a single value from 3 to 8 on every shape, thus getting “guessing correctly” on a specific category while scoring 0% on all others. These results demonstrate that models do not implement the $s_{1}+s_{2}-2$ algorithm.
@@ 3.4 Visually-Cued CoT
Table 4: Performance on heptagons and arrow on plus for Qwen2-VL, Molmo, Janus-Pro, and GPT models. Results are presented as direct / VC-CoT. When there are no image annotations, we use a standard CoT prompt denoted by $\dagger$. VC-CoT proves most effective in engaging MLLMs to count sides, compared to direct prompting.
@@ Appendix D What does the vision-encoder see?
LLaVA-1.5 0.95 0.95 0.65 0.50 0.44 0.48
LLaVA-Next 0.99 0.97 0.66 0.58 0.50 0.46
Qwen2-VL 1.00 0.98 0.84 0.67 0.51 0.50
InternVL 0.97 0.90 0.58 0.50 0.47 0.47
LLaVA-OneVision 1.00 1.00 0.82 0.58 0.52 0.52
LLaMA-3.2 0.87 0.79 0.40 0.37 0.36 0.34
Math-PUMA 1.00 1.00 75.50 56.82 50.71 50.56
G-LLaVA 0.96 0.97 0.65 0.48 0.44 0.47
Molmo 0.99 0.97 0.76 0.54 0.51 0.46
Janus-Pro 0.98 0.95 0.74 0.47 0.46 0.43
Math-LLaVA 0.96 0.97 0.65 0.48 0.44 0.48
@@ Appendix G Visually-Cued Chain-of-Thought Shape Results
Tables 10 and 11 present results for the remaining models not included in Table 4.

### v4

@@ Abstract
William Rudman⋆1, Michal Golovanevsky⋆1, Amir Bar2, Vedant Palit3, Yann LeCun4, Carsten Eickhoff5, Ritambhara Singh1 1Brown University, 2Tel Aviv University, 3IIT Kharagpur, 4New York University, 5University of Tübingen {william_rudman, michal_golovanevsky}@brown.edu
@@ 3.1 Probing Geometric Knowledge
3.1.1 Which visual features affect accuracy?
To better understand the causes of model errors, we analyzed the relationship between accuracy and five visual attributes: background color, shape color, size, rotation, and contrast ratio. The contrast ratio is defined as the difference between the background and shape colors. Table 2 shows the correlation coefficients between each factor and model accuracy for thirteen different models. A positive correlation means that accuracy improves as the factor increases, while a negative correlation suggests the factor has a harmful effect on accuracy.
Model Background Color Shape Color Size Rotation Contrast Ratio
LLaVA 1.5 -0.005 0.066 0.126 0.089 0.047
LLaVA 1.6 -0.197 -0.015 0.039 0.010 -0.110
Qwen-VL 0.127 0.009 0.096 -0.226 0.062
InternVL -0.042 0.061 0.186 0.011 -0.065
LLaVA-OneVision 0.012 0.033 0.026 -0.001 -0.020
LLaMA 3.2 0.021 0.026 0.049 0.010 0.014
Molmo 0.003 0.020 0.039 -0.270 -0.005
Janus -0.184 0.008 0.223 -0.144 -0.146
GPT-4-Turbo -0.056 0.051 0.035 -0.092 -0.023
GPT-4o 0.031 0.039 -0.054 -0.144 0.068
Math-LLaVA 0.122 -0.001 -0.006 -0.362 0.066
G-LLaVA 0.209 0.060 0.023 -0.040 0.137
Math-PUMA 0.018 -0.017 0.172 -0.079 -0.020
Abs Val Avg 0.079 0.031 0.083 0.098 0.060
Table 2: Correlation between visual features and model accuracy. Values represent how strongly each factor relates to correct predictions. Rotation and size have the highest average influence.
The results show that no single visual factor explains performance across all models. For example, rotation has a strong negative effect in Qwen2-VL and Math-LLaVA but has almost no effect in LLaVA-OneVision. On average, rotation shows the strongest overall influence, followed by size.
Contrast ratio, a factor that is often important in traditional vision tasks, shows mixed results. GPT-4o and G-LLaVA perform slightly better with higher contrast, while LLaVA 1.6 and Janus perform worse. These inconsistencies suggest that visual sensitivity varies significantly depending on model architecture and training.
Taken together, these findings suggest that misclassifications are not caused by a single weak point in the visual input. Instead, each model exhibits its own unique set of sensitivities. These insights complement our broader evaluation of shape recognition and motivate a closer look at the components of the system. In the next sections, we analyze the language model and the vision encoder independently to understand their contributions to these errors.
@@ 3.1.2 What does the text-decoder know?
  Name the <n >-sided polygon / Num of sides in <shape >
LM Only Accuracy (%)
LLaVA-1.5 67% / 100%
LLaVA-Next 100% / 100%
LLaVA-OneVision 100% / 100%
Qwen2-VL 100% / 100%
LLaMA-3.2 100% / 100%
InternVL 100% / 100%
Molmo 100% / 100%
Janus-Pro 100% / 100%
Math-LLaVA 100% / 100%
G-LLaVA 50% / 100%
Math-PUMA 100% / 100%
GPT-4o 100% / 100%
GPT-4-Turbo 100% / 100%
[…]
To assess LLMs’ knowledge of the geometric properties MLLMs struggle with, we construct the following prompts: (1) “What is the name of a <n>-sided polygon?” where <n> ranges from 3 to 8, and (2) “How many sides does a <shape> have?” where <shape> is a regular polygon from triangle to octagon. The results, shown in Table 9, reveal that the underlying language models know geometric properties. All models get 100% accuracy on “How many sides does a <shape> have?”, and 11 out of 13 models get 100% accuracy on “What is the name of a <n>-sided polygon?”. The success of LLM backbones in these tasks aligns with prior findings that state-of-the-art models perform well on lower-level math problems Xiang et al. (2025). Note that LLaVA 1.5 and G-LLaVA’s LMs struggle with naming certain shapes, suggesting some failures stem from gaps in pretraining data rather than vision-related issues. In Appendix E, we analyze Google N-Grams data and confirm that the shapes these models struggle with are also the least frequently mentioned in text.
@@ 3.3 Counting Sides of Abstract Shapes
Merged Shapes: We combine two regular polygons, such as a triangle and a square, where the shapes share a side. A simple shared-side algorithm can be applied to obtain the number of the sides: $s_{1}+s_{2}-2$ where $s_{1}$ and $s_{2}$ are the number of sides of shape 1 and shape 2, respectively. Applying this algorithm to a triangle and a square gives: 3+42=53+4-2=5 total sides. Table 5 shows that models generally fail to implement this algorithm, with most achieving 0% accuracy on merged shapes (Tri-Sq through Sq-Hep). LLaVA-OneVision and LLaMA-3.2 each achieve over 90% accuracy on exactly one merged shape. This is because most open-source models (except for Molmo, which is trained on counting) tend to guess a single value from 3 to 8 on every shape, thus getting “guessing correctly” on a specific category while scoring 0% on all others. These results demonstrate that models do not implement the $s_{1}+s_{2}-2$ algorithm.
@@ 3.4 Visually-Cued CoT
Table 6: Performance on heptagons and arrow on plus for Qwen2-VL, Molmo, Janus-Pro, and GPT models. Results are presented as direct / VC-CoT. When there are no image annotations, we use a standard CoT prompt denoted by $\dagger$. VC-CoT is more effective than direct CoT prompting for MLLMs in side counting.
@@ Appendix D What does the vision-encoder see?
LLaVA-1.5 95% 95% 65% 50% 44% 48%
LLaVA-Next 99% 97% 66% 58% 50% 46%
Qwen2-VL 100% 98% 84% 67% 51% 50%
InternVL 97% 90% 58% 50% 47% 47%
LLaVA-OneVision 100% 100% 82% 58% 52% 52%
LLaMA-3.2 87% 79% 40% 37% 36% 34%
Math-PUMA 100% 100% 76% 57% 51% 51%
G-LLaVA 96% 97% 65% 48% 44% 47%
Molmo 99% 97% 76% 54% 51% 46%
Janus-Pro 98% 95% 74% 47% 46% 43%
Math-LLaVA 96% 97% 65% 48% 44% 48%
@@ Appendix G Visually-Cued Chain-of-Thought Shape Results
Tables 12 and 13 present results for the remaining models not included in Table 6.